OpenStudy (anonymous):

Time for integration: x*sqrt(x^2 + 9)

OpenStudy (psymon):

You would just let u = x^2 + 9 u = x^2 + 9, so du = 2xdx and dx = du/2x

OpenStudy (anonymous):

\[\int\limits_{}^{}x \sqrt{x^{2}+9} dx\]

OpenStudy (psymon):

Does the u-substitution I mentioned make sense?

OpenStudy (anonymous):

thats where i started and then it went off the rails...i'll try again...brb

OpenStudy (psymon):

Alrighty.

OpenStudy (anonymous):

argh I was doing it right and then chickened out....and with confidence 2/3 * 1/2 will get to 1/3 which is in the answer thanks

OpenStudy (psymon):

Good job, np :3